OpenStudy (anonymous):

How do you solve this? 5 = log3(x^2 + 18)

OpenStudy (anonymous):

5 = log 10^5 --> log 10^5 = log 3 (x^2 +18). take both log out 10^5 = 3 (x^2 +18) and then solve as usual

OpenStudy (anonymous):

Awesome, thank you!

OpenStudy (anonymous):

np :)

OpenStudy (raden):

the 3 is as the base or no ?

OpenStudy (anonymous):

Yeah. It just says solve for X

OpenStudy (raden):

hmmm... is there a typo ? mabe like this :